Highlands Company uses the weighted-average method in its process costing system. It processes wood pulp for various manufacturers of paper products. Data relating to tons of pulp processed during June are provided below:  

Percent Completed
Tons of Pulp Materials Labor and Overhead
Work in process, June 1 80,400 89 % 23 %
Work in process, June 30 50,800 47 % 15 %
Started into production during June 301,800

Required:  

1. Compute the number of tons of pulp completed and transferred out during June.

2. Compute the equivalent units of production for materials and for labor and overhead for June.
Answer #1

Tons of pulp completed and transferred

= Beginning wip + Started into production - Ending wip

= 80,400 + 301,800 - 50,800

= 331,400

Equivalent units of production for materials

= 331,400 + (50,800*47%)

= 355,276

Equivalent units of production for labor

= 331,400 + (50,800*15%)

= 339,020